how many grams of lead 2 sulfate would be produced from the complete reaction of 23.6 grams of lead 2 oxide
1
Expert's answer
2016-04-26T12:17:06-0400
Lead (II) sulphate can be prepared by lead(II) oxide by warm sulphuric acid according to the following reaction:
23.6g xg
PbO+H2SO4=PbSO4+H2O
1m 1m
According to the stechiometric equation 1 mole of PbSO4 is produced from 1 mole of the PbO, so the amount of the PbSO4 will be calculated from the proportion x=(23,6×1)÷1=23.6g
x=23.6g